# Computing Balanced Solutions for Large International Kidney Exchange Schemes When Cycle Length Is Unbounded

### v1

@@ Abstract
In kidney exchange programmes (KEP) patients may swap their incompatible donors leading to cycles of kidney transplants. Nowadays, countries try to merge their national patient-donor pools leading to international KEPs (IKEPs). As shown in the literature, long-term stability of an IKEP can be achieved through a credit-based system. In each round, every country is prescribed a “fair” initial allocation of kidney transplants. The initial allocation, which we obtain by using solution concepts from cooperative game theory, is adjusted by incorporating credits from the previous round, yielding the target allocation. The goal is to find, in each round, an optimal solution that closely approximates this target allocation. There is a known polynomial-time algorithm for finding an optimal solution that lexicographically minimizes the country deviations from the target allocation if only $2$-cycles (matchings) are permitted. In practice, kidney swaps along longer cycles may be performed. However, the problem of computing optimal solutions for maximum cycle length $\ell$ is NP-hard for every $\ell\geq 3$. This situation changes back to polynomial time once we allow unbounded cycle length. However, in contrast to the case where $\ell=2$, we show that for $\ell=\infty$, lexicographical minimization is only polynomial-time solvable under additional conditions (assuming ${\sf P}\neq{\sf NP}$). Nevertheless, the fact that the optimal solutions themselves can be computed in polynomial time if $\ell=\infty$ still enables us to perform a large scale experimental study for showing how stability and total social welfare are affected when we set $\ell=\infty$ instead of $\ell=2$. omputational complexity, cooperative game theory, partitioned permutation game, international kidney exchange
@@ 1 Introduction
In this paper, we introduce a new class of cooperative games called partitioned permutation games, which are closely related to the known classes of permutation games and partitioned matching games (introduced at AAMAS 2019 [13]). Both partitioned matching games and partitioned permutation games have immediate applications in international kidney exchange. Before defining these games, we first explain this application area.
@@
Kidney Exchange. The most effective treatment for kidney failure is transplanting a kidney from a deceased or living donor, with better long-term outcomes in the latter case. However, a kidney from a family member or friend might be medically incompatible and could easily be rejected by the patient’s body. Therefore, many countries are running national Kidney Exchange Programmes (KEPs) [12]. In a KEP, all patient-donor pairs are placed together in one pool. If for two patient-donor pairs $(p,d)$ and $(p^{\prime},d^{\prime})$, it holds that $d$ and $p$ are incompatible, as well as $d^{\prime}$ with $p^{\prime}$, then $d$ and $d^{\prime}$ could donate a kidney to $p^{\prime}$ and $p$, respectively. This is a $2$-way exchange.
@@
We now generalize a $2$-way exchange. We model a pool of patient-donor pairs as a directed graph $G=(V,A)$ (the compatibility graph) in which $V$ consists of the patient-donor pairs, and $A$ consists of every arc $(u,v)$ such that the donor of $u$ is compatible with the patient of $v$. In a directed cycle $C=u_{1}u_{2}\ldots u_{k}u_{1}$, for some $k\geq 2$, the kidney of the donor of $u_{i}$ could be given to the patient of $u_{i+1}$ for every $i\in\{1,\ldots,k+1\}$, with $u_{k+1}:=u_{1}$. This is a $k$-way exchange using the exchange cycle $C$. To prevent exchange cycles from breaking (and a patient from losing their willing donor), hospitals perform the $k$ transplants in a $k$-way exchange simultaneously. Hence, KEPs impose a bound $\ell$ (the exchange bound) on the maximum length of an exchange cycle, typically $2\leq\ell\leq 5$. An $\ell$-cycle packing of $G$ is a set ${\mathcal{C}}$ of directed cycles, each of length at most $\ell$, that are pairwise vertex-disjoint; if $\ell=\infty$, we also say that ${\mathcal{C}}$ is a cycle packing. The size of ${\mathcal{C}}$ is the number of arcs that belong to a cycle in ${\mathcal{C}}$.
@@
KEPs operate in rounds. A solution for round $r$ is an $\ell$-cycle packing in the corresponding compatibility graph $G^{r}$. The goal is to help as many patients as possible in each round. Hence, to maximize the number of transplants in round $r$, we seek for an optimal solution, that is, a maximum $\ell$-cycle packing of $G^{r}$, i.e., one that has maximum size. After a round, some patients have received a kidney or died, and other patient-donor pairs may have arrived, resulting in a new compatibility graph $G^{r+1}$ for round $r+1$.
@@
The main computational issue for KEPs is how to find an optimal solution in each round. If $\ell=2$, we can transform, by keeping the “double” arcs, a computability graph $G$ into an undirected graph $D=(V,E)$ as follows. For every $u,v\in V$, we have $uv\in E$ if and only if $(u,v)\in A$ and $(v,u)\in A$. It then remains to compute, in polynomial time, a maximum matching in $D$. If we set $\ell=\infty$, a well-known trick works (see e.g. [1]). We transform $G$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $u\in V$ and its copy $u^{\prime}\in V^{\prime}$, we add the edge $uu^{\prime}$ with weight 00. For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. Now it remains to find in polynomial time a maximum weight perfect matching in $H$. However, for any constant $\ell\geq 3$, the situation changes, as shown by Abraham, Blum and Sandholm [1].
@@
International Kidney Exchange. As merging pools of national KEPs leads to better outcomes, the focus nowadays lies on forming an international KEP (IKEP), e.g. Austria and the Czech Republic [17]; Denmark, Norway and Sweden; and Italy, Portugal and Spain [41]. Apart from ethical, legal and logistic issues (all beyond our scope), there is now a new and highly non-trivial issue that needs to be addressed: How can we ensure long-term stability of an IKEP? If countries are not treated fairly, they may leave the IKEP, and it could even happen that each country runs their own national KEP again.
Example. Let $G$ be the compatibility graph from Figure 1. Then a total of five kidney transplants is possible if the exchange cycle $C=abdeca$ is used. In that case, countries $1$, $2$ and $3$ receive three, two and zero kidney transplants, respectively. ∎
@@
Cooperative Game Theory considers fair distributions of common profit if all parties involved collaborate with each other. Before describing its role in our setting, we first give some terminology. A (cooperative) game is a pair $(N,v)$, where $N$ is a set of $n$ players and $v:2^{N}\to\mathbb{R}$ is a value function with $v(\emptyset)=0$. A subset $S\subseteq N$ is a coalition. If for every possible partition $(S_{1},\ldots,S_{r})$ of $N$ it holds that $v(N)\geq v(S_{1})+\cdots+v(S_{r})$, then players will benefit most by forming the grand coalition $N$. The problem is then how to fairly distribute $v(N)$ amongst the players of $N$. An allocation is a vector $x\in\mathbb{R}^{N}$ with $x(N)=v(N)$ (we write $x(S)=\sum_{p\in S}x_{p}$ for $S\subseteq N$). A solution concept prescribes a set of fair allocations for a game $(N,v)$. Here, the notion of fairness depends on context.
The solution concepts we consider are the Banzhaf value, Shapley value, nucleolus, benefit value, contribution value (see Section 4) and the core. They all prescribe a unique allocation except for the core, which consists of all allocations $x\in\mathbb{R}^{N}$ with $x(S)\geq v(S)$ for every $S\subseteq N$. Core allocations ensure $N$ is stable, as no subset $S$ will benefit from forming their own coalition. But a core may be empty.
We now define some relevant games. For a directed graph $G=(V,A)$ and subset $S\subseteq V$, we let $G[S]=(S,\{(u,v)\in A\;|\;u,v\in S\})$ be the subgraph of $G$ induced by $S$. An $\ell$-permutation game on a graph $G=(V,E)$ is the game $(N,v)$, where $N=V$ and for $S\subseteq N$, the value $v(S)$ is the maximum size of an $\ell$-cycle packing of $G[S]$. Two special cases are well studied. We obtain a matching game if $\ell=2$, which may have an empty core (e.g. when $G$ is a triangle) and a permutation game if $\ell=\infty$, whose core is always nonempty [39].
@@
To ensure IKEP stability, we use the model of Klimentova et al. [31], which is a credit-based system. For round $r\geq 1$, let $G^{r}$ be the computability graph with associated game $(N^{r},v^{r})$; let $y^{r}$ be the initial allocation (as prescribed by some solution concept ${\mathcal{S}}$); and let $c^{r}:N^{r}\to\mathbb{R}$ be a credit function, which satisfies $\sum_{p\in N{{}^{r}}}c^{r}_{p}=0$; if $r=1$, we set $c^{r}\equiv 0$. For $p\in N$, we set $x^{r}_{p}:=y^{r}_{p}+c^{r}_{p}$ to obtain the target allocation $x^{r}$ for round $r$ (which is indeed an allocation, as $y^{r}$ is an allocation and $\sum_{p\in N}c_{p}^{r}=0$). We choose some maximum $\ell$-cycle packing ${\mathcal{C}}$ of $G^{r}$ as optimal solution for round $r$ (out of possibly exponentially many optimal solutions). Let $s_{p}({\mathcal{C}})$ be the number of kidney transplants for patients in country $p$ (with donors both from $p$ and other countries). For $p\in N$, we set $c^{r+1}_{p}:=x^{r}_{p}-s_{p}({\mathcal{C}})$ to get the credit function $c^{r+1}$ for round $r+1$ (note that $\sum_{p\in N}c_{p}^{r+1}=0$). For round $r+1$, a new initial allocation $y^{r+1}$ is prescribed by ${\mathcal{S}}$ for the associated game $(N^{r+1},v^{r+1})$. For every $p\in N$, we set $x_{p}^{r+1}:=y_{p}^{r+1}+c_{p}^{r+1}$, and we repeat the process.
@@
Related Work. Benedek et al. [8] proved the following theorem for $\ell=2$ (the “matching” case); in contrast, Biró et al. [13] showed that it is NP-hard to find a weakly close maximum matching even for $|N|=2$ once the games are defined on edge-weighted graphs.
@@
Benedek et al. [8] used the algorithm of Theorem 1.2 to perform simulations for up to fifteen countries for $\ell=2$. As initial allocations, they used the Shapley value, nucleolus, benefit value and contribution value, with the Shapley value yielding the best results (together with the Banzhaf value in the full version [9] of [8]). This is in line with the results of Klimentova et al. [31] and Biró et al. [11] for $\ell=3$. Due to Theorem 1.1, the simulations in [11, 31] are for up to four countries and use weakly close optimal solutions.
For $\ell=2$, we refer to [38] for an alternative model based on so-called selection ratios using lower and upper target numbers. IKEPs have also been modelled as non-cooperative games in the consecutive matching setting, which has $2$-phase rounds: national pools in phase 1 and a merged pool for unmatched patient-donor pairs in phase 2; see [19, 18, 37] for some results in this setting.
Fairness (versus optimality) issues are also studied for national KEPs, in particular in the US, which is different from Europe (our setting). Namely, hospitals in the US are more independent and are given positive credits to register their easy-to-match patient-donor pairs to one of the three national KEPs, and negative credits for registering their hard-to-match pairs. The US setting is extensively studied [3, 4, 2, 16, 27, 40], but beyond the scope of our paper.
Our Results. We consider partitioned permutation games and IKEPs, so we assume $\ell=\infty$. This assumption is not realistic in kidney exchange, but has also been made for national KEPs to obtain general results of theoretical interest [5, 14, 15, 34] and may have wider applications (e.g. in financial clearing). We also aim to research how stability and total number of kidney transplants are affected when moving from one extreme ($\ell=2$) to the other ($\ell=\infty$). As such our paper consists of a theoretical and experimental part.
@@
The core of every partitioned permutation game is non-empty, and it is possible to find a core allocation in polynomial time. Moreover, for partitioned permutation games of fixed width $c$, the problem of deciding if an allocation is in the core is polynomial-time solvable if $c=1$ and coNP-complete if $c\geq 2$.
Due to Theorem 1.1, we cannot hope to generalize Theorem 1.2 to hold for any constant $\ell\geq 3$. Nevertheless, Theorem 1.1 leaves open the question if Theorem 1.2 is true for $\ell=\infty$ (the “cycle packing” case) instead of only for $\ell=2$ (the “matching” case). We show that the answer to this question is no (assuming ${\sf P}\neq{\sf NP}$).
@@
Our theoretical results which highlighted severe computational limitations, and we now turn to our simulations. These are guided by our theoretical results. Namely, we note that the algorithm in Theorem 1.5 is neither practical for instances of realistic size (which we aim for) nor acceptable in the setting of kidney exchange due to being a randomized algorithm. Therefore and also due to Theorem 1.4, we formulate the problems of computing a weakly or strongly close optimal solution as integer linear programs, as described in Section 3. This enables us to use an ILP solver. We still exploit the fact that for $\ell=\infty$ (Theorem 1.1) we can find optimal solutions and values $v(S)$ in polynomial time. In this way we can perform, in Section 4, simulations for IKEPs up to ten countries, so more than the four countries in the simulations for $\ell=3$ [11, 31], but less than the fifteen countries in the simulations for $\ell=2$ [8].
For the initial allocations we use two easy-to-compute solution concepts: the benefit value and contribution value, and three hard-to-compute solution concepts: the Banzhaf value, Shapley value, and nucleolus. Our simulations show, just like those for [8], that a credit system using strongly close optimal solutions makes an IKEP the most balanced without decreasing the overall number of transplants. The exact improvement is determined by the choice of solution concept. Our simulations indicate that the Banzhaf value yields the best results: on average, a deviation of up to 0.90% from the target allocation. Moreover, moving from $\ell=2$ to $\ell=\infty$ yields on average 46% more kidney transplants (using the same simulation instances generated by the data generator [33]). However, the exchange cycles may be very large, in particular in the starting round.

### v2

# Márton Benedek 1122    Péter Biró 2211    Gergely Csáji 22 Matthew Johnson 33

@@ Abstract
In kidney exchange programmes (KEP) patients may swap their incompatible donors leading to cycles of kidney transplants. Recently, countries try to merge their national patient-donor pools leading to international KEPs (IKEPs). As shown in the literature, long-term stability of an IKEP can be achieved through a credit-based system. In each round, every country is prescribed a “fair” initial allocation of kidney transplants. The initial allocation, which we obtain by using solution concepts from cooperative game theory, is adjusted by incorporating credits from the previous round, yielding the target allocation. The goal is to find, in each round, an optimal solution that closely approximates this target allocation. There is a known polynomial-time algorithm for finding an optimal solution that lexicographically minimizes the country deviations from the target allocation if only $2$-cycles (matchings) are permitted. In practice, kidney swaps along longer cycles may be performed. However, the problem of computing optimal solutions for maximum cycle length $\ell$ is NP-hard for every $\ell\geq 3$. By contrast, the problem is polynomial time solvable once we allow unbounded cycle length. However, in contrast to the case where $\ell=2$, we show that for $\ell=\infty$, lexicographical minimization is only polynomial-time solvable under additional conditions (assuming ${\sf P}\neq{\sf NP}$). Nevertheless, the fact that the optimal solutions themselves can be computed in polynomial time if $\ell=\infty$ still enables us to perform a large scale experimental study for showing how stability and total social welfare are affected when we set $\ell=\infty$ instead of $\ell=2$. omputational complexity, cooperative game theory, partitioned permutation game, international kidney exchange
@@ 1 Introduction
We introduce a new class of cooperative games called partitioned permutation games, which are closely related to the known classes of permutation games [40] and partitioned matching games [14]. Partitioned matching games and partitioned permutation games have immediate applications in international kidney exchange. Before defining these games, we first explain this application area.
@@
Kidney Exchange. The most effective treatment for kidney failure is transplanting a kidney from a deceased or living donor, with better long-term outcomes in the latter case. However, a kidney from a family member or friend might be medically incompatible and could easily be rejected by the patient’s body. Therefore, many countries run national Kidney Exchange Programmes (KEPs) [13]. In a KEP, all patient-donor pairs are placed together in one pool. If for two patient-donor pairs $(p,d)$ and $(p^{\prime},d^{\prime})$, it holds that $d$ and $p$ are incompatible, as well as $d^{\prime}$ with $p^{\prime}$, then $d$ could donate a kidney to $p^{\prime}$, and $d^{\prime}$ could donate a kidney to $p$. This is a $2$-way exchange.
@@
We now generalize a $2$-way exchange. We model a pool of patient-donor pairs as a directed graph $G=(V,A)$ (the compatibility graph) in which each vertex of $V$ is a patient-donor pair, and $A$ consists of every arc $(u,v)$ such that the donor of $u$ is compatible with the patient of $v$. In a directed cycle $C=u_{1}u_{2}\ldots u_{k}u_{1}$, for some $k\geq 2$, the kidney of the donor of $u_{i}$ could be given to the patient of $u_{i+1}$ for every $i\in\{1,\ldots,k+1\}$, with $u_{k+1}:=u_{1}$. This is a $k$-way exchange using the exchange cycle $C$. To prevent exchange cycles from breaking (and a patient from losing their willing donor), hospitals perform the $k$ transplants in a $k$-way exchange simultaneously. Hence, KEPs impose a bound $\ell$, called the exchange bound, on the maximum length (number of edges) of an exchange cycle, typically $2\leq\ell\leq 5$. An $\ell$-cycle packing of $G$ is a set ${\mathcal{C}}$ of directed cycles, each of length at most $\ell$, that are pairwise vertex-disjoint; if $\ell=\infty$, we also say that ${\mathcal{C}}$ is a cycle packing. The size of ${\mathcal{C}}$ is the number of arcs that belong to a cycle in ${\mathcal{C}}$, or equivalently, the sum of the lengths of the cycles of ${\cal C}$.
@@
KEPs operate in rounds. A solution for round $r$ is an $\ell$-cycle packing in the corresponding compatibility graph $G^{r}$. The goal is to help as many patients as possible in each round. Hence, to maximize the number of transplants in round $r$, we seek an optimal solution, that is, a maximum $\ell$-cycle packing of $G^{r}$, i.e., one that has maximum size. After round $r$, some patients will have received a kidney or died, while other patient-donor pairs may have arrived. This results in a new compatibility graph $G^{r+1}$ for round $r+1$.
@@
The main computational issue for KEPs is how to find an optimal solution in each round. If $\ell=2$, we can transform, by keeping the “double” arcs, a compatibility graph $G$ into an undirected graph $D=(V,E)$ as follows. For every $u,v\in V$, we have $uv\in E$ if and only if $(u,v)\in A$ and $(v,u)\in A$. It then remains to compute a maximum matching in $D$, which can be done in polynomial time. If we set $\ell=\infty$, a well-known trick works (see e.g. [1]). We transform $G$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $u\in V$ and its copy $u^{\prime}\in V^{\prime}$, we add the edge $uu^{\prime}$ with weight 00. For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. Now it remains to find in polynomial time a maximum weight perfect matching in $H$. However, for any constant $\ell\geq 3$, the complexity changes, as shown by Abraham, Blum and Sandholm [1].
@@
International Kidney Exchange. As merging pools of national KEPs leads to better outcomes, many countries today work together with the aim of forming an international KEP (IKEP), e.g. Austria and the Czech Republic [18]; Denmark, Norway and Sweden; and Italy, Portugal and Spain [43]. Apart from ethical, legal and logistical issues (all beyond our scope), there is now a new and highly non-trivial issue that needs to be addressed: How can we ensure long-term stability of an IKEP? If countries are not treated fairly, they may leave the IKEP. In a worst-case scenario it could even happen that in the end each country returns to their own national KEP again.
Example. Let $G$ be the compatibility graph from Figure 1. Then a total of five kidney transplants is possible (only) if we choose the solution of size $5$ that consists of the single exchange cycle $C=abdeca$. We have that patient-donor pairs $a,b,c$ belong to country $1$, whereas patient-donor pairs $d$ and $e$ belong to country $2$, and patient-donor pair $f$ belongs to country $3$. So using this solution, countries $1$, $2$ and $3$ receive three, two and zero kidney transplants, respectively. Note that without cooperation only country $2$ can receive kidney transplants in this example.∎
@@
Cooperative Game Theory considers fair distributions of joint profit if all parties involved collaborate (the notion of fairness depends on context). Before describing its role in our setting, we first give some terminology. A (cooperative) game is a pair $(N,v)$, where $N$ is a set of $n$ players and $v:2^{N}\to\mathbb{R}$ is a value function with $v(\emptyset)=0$. A subset $S\subseteq N$ is a coalition. If for every possible partition $(S_{1},\ldots,S_{r})$ of $N$ it holds that $v(N)\geq v(S_{1})+\cdots+v(S_{r})$, then players will benefit most by forming the grand coalition $N$. The problem is then how to fairly distribute $v(N)$ amongst the players of $N$. An allocation is a vector $x\in\mathbb{R}^{N}$ with $x(N)=v(N)$ (we write $x(S)=\sum_{p\in S}x_{p}$ for $S\subseteq N$). A solution concept prescribes a set of fair allocations for a game $(N,v)$.
The solution concepts considered in this paper are the Shapley value, Banzhaf value (two variants) nucleolus, tau value, benefit value, contribution value (all of which are defined in Section 4) and the core. They all prescribe a unique allocation except for the core, which consists of all allocations $x\in\mathbb{R}^{N}$ with $x(S)\geq v(S)$ for every $S\subseteq N$. Core allocations ensure $N$ is stable, as no subset $S$ will benefit from forming their own coalition. But a core may be empty.
We now define some relevant games. For a directed graph $G=(V,A)$ and subset $S\subseteq V$, we let $G[S]=(S,\{(u,v)\in A\;|\;u,v\in S\})$ be the subgraph of $G$ induced by $S$. An $\ell$-permutation game on a directed graph $G=(V,A)$ is the game $(N,v)$, where $N=V$ and for $S\subseteq N$, the value $v(S)$ is the maximum size of an $\ell$-cycle packing of $G[S]$. Two special cases are well studied. We obtain a matching game if $\ell=2$, which may have an empty core (e.g. when $G$ is a triangle on vertices $a,b,c$ with arcs $(a,b)$, $(b,a)$, $(a,c)$, $(c,a)$, $(b,c)$, $(c,b)$) and a permutation game if $\ell=\infty$, whose core is always nonempty [40].
@@
To ensure IKEP stability, we use the model of Klimentova et al. [32], which is a credit-based system. For round $r\geq 1$, let $G^{r}$ be the compatibility graph with associated game $(N^{r},v^{r})$; let $y^{r}$ be the initial allocation (as prescribed by some solution concept ${\mathcal{S}}$); and let $c^{r}:N^{r}\to\mathbb{R}$ be a credit function, which satisfies $\sum_{p\in N{{}^{r}}}c^{r}_{p}=0$; if $r=1$, we set $c^{r}\equiv 0$. For $p\in N$, we set $x^{r}_{p}:=y^{r}_{p}+c^{r}_{p}$ to obtain the target allocation $x^{r}$ for round $r$; note that $x^{r}$ is indeed an allocation, as $y^{r}$ is an allocation and $\sum_{p\in N}c_{p}^{r}=0$). We choose some maximum $\ell$-cycle packing ${\mathcal{C}}$ of $G^{r}$ as optimal solution for round $r$ (out of possibly exponentially many optimal solutions). Let $s_{p}({\mathcal{C}})$ be the number of kidney transplants for patients in country $p$ (with donors both from $p$ and other countries). For $p\in N$, we set $c^{r+1}_{p}:=x^{r}_{p}-s_{p}({\mathcal{C}})$ to get the credit function $c^{r+1}$ for round $r+1$ (note that $\sum_{p\in N}c_{p}^{r+1}=0$). For round $r+1$, a new initial allocation $y^{r+1}$ is prescribed by ${\mathcal{S}}$ for the associated game $(N^{r+1},v^{r+1})$. For every $p\in N$, we set $x_{p}^{r+1}:=y_{p}^{r+1}+c_{p}^{r+1}$, and we repeat the process.
@@
Related Work
Benedek et al. [9] proved the following theorem for $\ell=2$ (the “matching” case); in contrast, Biró et al. [14] showed that it is NP-hard to find a weakly close maximum matching even for $|N|=2$ once the games are defined on edge-weighted graphs.
@@
Benedek et al. [9] used the algorithm of Theorem 1.2 to perform simulations for up to 15 countries for $\ell=2$. As initial allocations, they used the Shapley value, nucleolus, benefit value and contribution value, with the Shapley value yielding the best results. Afterwards, Benedek et al. [10] extended the experimental results from [9] for $\ell=2$ by also considering two variants of the Banzhaf value and the tau value. It turned out that the Banzhaf value variants behaved slightly better than the Shapley value. The good performance of the Shapley value in [9, 10] is in line with the simulation results of Klimentova et al. [32] and Biró et al. [12] for $\ell=3$. Due to Theorem 1.1, the simulations in [12, 32] are for up to four countries and use weakly close optimal solutions.
For $\ell=2$, we refer to [39] for an alternative model based on so-called selection ratios using lower and upper target numbers. IKEPs have also been modelled as non-cooperative games in the consecutive matching setting. In this setting, each round consists of two phases: national pools in phase 1 and a merged pool for unmatched patient-donor pairs in phase 2; see [19, 20, 38] for some results in this setting.
Fairness (versus optimality) issues are also studied for national KEPs, in particular in the US. The US is different from Europe, the setting we consider, as hospitals in the US are more independent and are given positive credits to register their easy-to-match patient-donor pairs to one of the three national KEPs, and negative credits for registering their hard-to-match pairs. The US setting is extensively studied [2, 3, 4, 17, 28, 42], but beyond the scope of our paper.
Our Results
We consider partitioned permutation games and IKEPs, so we assume $\ell=\infty$. This assumption is not realistic in kidney exchange, but has also been made for national KEPs to obtain general results of theoretical interest [5, 15, 16, 35] and may have wider applications (e.g. in financial clearing). We also aim to research how stability and the total number of kidney transplants are affected when moving from one extreme ($\ell=2$) to the other ($\ell=\infty$). As such our paper consists of a theoretical part and an experimental part.
@@
The core of every partitioned permutation game is nonempty, and it is possible to find a core allocation in polynomial time. Moreover, for partitioned permutation games of fixed width $c$, the problem of deciding if an allocation is in the core is polynomial-time solvable if $c=1$ and coNP-complete if $c\geq 2$.
Due to Theorem 1.1, we cannot hope to generalize Theorem 1.2 to hold for any constant $\ell\geq 3$. Nevertheless, Theorem 1.1 leaves open the question of whether there is an analogue of Theorem 1.2 for $\ell=\infty$, the “cycle packing” case; Theorem 1.2 is concerned only with $\ell=2$, the “matching” case. We show that the answer to this question is no (assuming ${\sf P}\neq{\sf NP}$).
@@
Our theoretical results highlighted severe computational limitations, and we now turn to our simulations. We perform a large-scale experimental study, in which our simulations are strongly guided by our theoretical results. Namely, we note that the algorithm in Theorem 1.5 is not practical for instances of realistic size (which we aim for). Moreover, it is a randomized algorithm. Hence, it it not acceptable either in the setting of kidney exchange, as policy makers would only use solution yielding the maximum number of kidney transplants. Therefore and also due to Theorem 1.4, we formulate the problems of computing a weakly or strongly close optimal solution as integer linear programs, as described in Section 3. This enables us to use an ILP solver. We still exploit the fact that for $\ell=\infty$ (Theorem 1.1) we can find optimal solutions and values $v(S)$ in polynomial time. In this way we can still perform, in Section 4, simulations for IKEPs up to ten countries, so more than the four countries in the simulations for $\ell=3$ [12, 32], but less than the 15 countries in the simulations for $\ell=2$ [10].
For the initial allocations we use two easy-to-compute solution concepts: the benefit value and contribution value and four hard-to-compute solution concepts: the Banzhaf value, Shapley value, nucleolus and tau-value. We also use the modified Banzhaf value called Banzhaf* value from [10], where the credits are incorporated directly into the coalitional values. As in [10], we research both the case where all countries have the same size and the case where countries can be either of small, medium or large size.
Our simulations show, as in [10], that a credit system using strongly close optimal solutions instead of weakly optimal solutions makes an IKEP up to 55% more balanced, without decreasing the overall number of transplants. The exact improvement is determined by the choice of solution concept. Our simulations indicate that the Banzhaf* value yields the best results, namely on average, a deviation of at most 0.90% from the initial allocation (but the differences between the Banzhaf* value with the Banzhaf value and Shapley value are small). Moreover, moving from $\ell=2$ to $\ell=\infty$ yields on average 46% more kidney transplants (using the same simulation instances generated by the data generator [34]). However, the exchange cycles may be very large, in particular in the starting round.